XYZ Corporation has just issued a callable (at Par) three year 7% coupon with semi-annual coupon payments. The bond can be called at par in two years or anytime thereafter on a coupon payment date. It has a price of $98. What is the bond's yield to maturity and yield call?

a. The yield to maturity is

B. what is the yield to call
Answer #1

For calculation of YTM, coupon rate is calculated as (100 * 0.07) / 2 = 3.5

No of periods = no of years * 2 = 3 * 2 = 6

From the excel calculations the yield has come out to be 3.88%. But it is semiannual yield.

Therefore annual yield to maturity = 3.88% * 2 = 7.76%

Now lets calculate yield to callHomeInsert Page Layout Formulas Data Review View Developer cut Copy Σ AutoSum Callbri Fill o-A- 困Merge & Center, $ , % , Md C Here yield to call is 4.05%. But this is semiannual rate.

Annual yield to call = 4.05% * 2 = 8.10%